the perimeter of a rectangle is 60 meters if its length exceeds its breath by 2 meter find its area
Answer by sachi(548)

let the breadth=x m
the the length = x+2 m
so the perimeter= 2(x+x+2)=60
or 2x+2=60/2=30
or 2x=30-2=28
or x=28/2=14=breadth
the length=14+2=16 m
so the area=14*16=224 m sqr